OpenStudy (anonymous):

find the average rate of change for the given function from x=1 to x=3 X Y -1 1 1 5 3 9

OpenStudy (hugsnotughs):

Welcome to Openstudy. :) Do you know the formula used to find the average rate of change?

OpenStudy (anonymous):

Yes I finally understand how to find the average rate of change. but I have another question regarding rate of change

OpenStudy (anonymous):

graph and find the average rate of change for the given function for x=1 to x+2 I do not understand how i would graph that if thats all the info they gave me help please?

OpenStudy (hugsnotughs):

\[Average~rate~of~change~from~x_{1}~\to~x_{2}=\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\]So...\[Average~rate~of~change~from~1~\to~3=\frac{ 9-5 }{ 3-1 }\] So just solve that and you will have your average rate of change. :) Now, what's ur question? c:

OpenStudy (anonymous):

I got 2 by the way (:

OpenStudy (anonymous):

graph and find the average rate of change for the given function for x=1 to x=2 I do not understand how i would graph that if thats all the info they gave me help please?

OpenStudy (hugsnotughs):

Hm...I am also confused...it is only saying what you need to find the Average Rate of Change of and not giving the actual graph or at least a graph that we should graph ourselves...

OpenStudy (anonymous):

Find the average rate of change for the given function from x = 1 to x = 2. (5 points) Select one: a. -2 b. c. d. 2

OpenStudy (hugsnotughs):

Oh okay, if that is the graph, and you have to find the average rate of change from x = 1 to x = 2. Great! Let's see... x1 = 1; x2 = 2; y1 = 3; y2 = 5 Let's solve this...:\[\frac{ 5-3 }{ 2-1 }=\frac{ 2 }{ 1 }=2\]

OpenStudy (anonymous):

did you get the points off the graph right does it matter what points??

OpenStudy (hugsnotughs):

Yes, it does. x=1 would be... (1, 3). This is hard to explain through text, but I hope you understand. :) They give you the x-value, but not the y-value. They want you to look back in the graph and find it.

OpenStudy (anonymous):

Oh okay I see what you mean omg thank you!!!! :D Can you help me with one more problem too? f(x)=2x+1 and g(x)=3x-1/2 solve for g(f((3)) what is this?????!

OpenStudy (hugsnotughs):

OMG, WHAT IS THAT?! (its a trick question) Okay, let's work this out...step by step. So it wants you to find g(f(3)). To do this, first you have to find f(3)... \[f(x)=2x+1\]\[f(3)=2(3)+1\]\[f(3)=6+1\]\[f(3)=7\] Now that we have f(3), we can find g(f(3))...which is now g(7). Do you think you can do the rest? All you have to do is fin g(7) now. :)

OpenStudy (anonymous):

I got 10? 3(7)-1/2 21-1/2=20/2 equals 10

OpenStudy (hugsnotughs):

Hm...you got it wrong. But I see it is because you didn't make the fraction into a numeral. Remember 1/2 = 0.5 ~So... it would be 21 - 0.5 = 20.5 ~I see where you are confused though... :)

OpenStudy (anonymous):

so its 20.5? yeah I'm confused lol

OpenStudy (hugsnotughs):

Yep. :)

OpenStudy (hugsnotughs):

Also, if you'll excuse me, I have to sleep, it's almost 4am where I'm at. >_< I hope I helped :)

OpenStudy (anonymous):

Thank you so much!!! (: